THE ROTATIONAL SPECTRA AND STRUCTURES OF THE HCCH-(OCS)_{2} and OCS-(HCCH)_{2} TRIMERS

\underline{SEAN A. PEEBLES} and ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, 930 North University Ave., Ann Arbor, MI 48109-1055 USA.

The two mixed trimers HCCH-(OCS)_{2} and OCS-(HCCH)_{2} have been identified and assigned by FTMW spectroscopy. Rotational constants and dipole moment components for both species are consistent with triangular, barrel-shaped configurations of the monomers.

The assignment of the rotational spectra of eight isotopomers for HCCH-(OCS)_{2} enabled an unambiguous structure determination and the discovery that this trimer contains a polar OCS dimer fragment. The two OCS molecules are found to be aligned in an unusual parallel arrangement with their dipole moments reinforcing one another.

Four isotopomers of the OCS-(HCCH)_{2} trimer were assigned. Due to a lack of isotopic data, a full structure determination was not possible. Semi-empirical calculations did, however, give a structure that aligned the acetylene monomers in a geometry intermediate between T-shaped and parallel that was consistent with the experimental data.